OpenStudy (anonymous):

how do you find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . .

OpenStudy (mertsj):

Is it arithmetic or geometric?

OpenStudy (anonymous):

@Mertsj it doesn't say :\

OpenStudy (mertsj):

I know. You have to figure it out.

OpenStudy (mertsj):

Do you always add the same number or multiply by the same number to get the next term.

OpenStudy (anonymous):

@Mertsj you add -5 so it's arithmetic?

OpenStudy (mertsj):

yes. and you have identified the common difference. So now hunt up your formula for the sum of an arithmetic sequence and plug into it.

OpenStudy (mertsj):

First you need to find the 12th term. Can you do that?

OpenStudy (mertsj):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

@Mertsj -53 is the 12th term

OpenStudy (mertsj):

Use that to find the 12th term.

OpenStudy (mertsj):

I got -54

OpenStudy (mertsj):

\[a _{12}=1+(12-1)(-5)=1-55=-54\]

OpenStudy (mertsj):

\[S _{12}=\frac{12(1-54)}{2}\]

OpenStudy (anonymous):

@Mertsj sorry thats what i was supposed to put up. do i plug -54 into Sn=n(a1+an)/2 as an?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

@Mertsj hen it's 318?

OpenStudy (mertsj):

I got -318

OpenStudy (anonymous):

thank you!!!! @Mertsj

OpenStudy (mertsj):

yw